# An exact dimension-reduced dynamic theory for developable surfaces and curve-fold origami

### v2

@@ Abstract
Curve-fold origami, composed of developable panels joined along a curved crease, exhibits rich dynamic behaviors relevant to metamaterials and soft robotic systems. Despite multiple approximated models (shell model, Kirchhoff rod model, etc.), a comprehensive and exact dynamical theory for curve-fold origami remains absent, limiting the precise prediction of its dynamics, especially for those with wide panels. In this work, we develop an exact dimension-reduced theory that focuses on the dynamics of curve-fold origami, utilizing the intrinsic one-dimensional nature of developable surfaces. As is well known in differential geometry, a curve on a developable surface determines the local frame and the shape of the developable surface under isometric constraint, making the problem indeed one-dimensional. Starting from a single developable surface, we investigate the kinematics and kinetic energy of a moving developable surface. By overcoming the difficulty of describing the motion of local frames, we derive the exact velocity field of wide surfaces solely described by the motion of the reference curve, which leads to the kinetic energy of the entire surface. The elastic energy follows the Wunderlich functional applicable to wide panels. Owing to the one-dimensional feature, the Lagrangian of the system, composed of both kinetic and elastic energy, is a functional of the reference curve. Thus, we may variate the Lagrangian and derive a nonlinear dynamical theory for the reference curve, which comprises governing equations similar to the rod model but can precisely describe the motion of developable surfaces. The theory is validated consistently in both Lagrangian and Eulerian frameworks and is further extended to curved-fold origami modeled as a coupled bi-rod system. Utilizing our exact 1D model, we theoretically analyze the dynamical behaviors of various developable structures, revealing that the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces, which are validated by finite element analysis quantitatively. Our exact dimensional-reduced dynamical theory captures the motions of developable surfaces and curve-fold origami (with wide panels) accurately at relatively low computational cost (since it is 1D), offering a potential contribution to the design of dynamical developable structures.
@@ 1 Introduction
Previous studies on the geometry and statics of curve-fold origami are under isometric assumptions, since the paper-like panels are unstretchable. Therefore, the panels will deform as developable surfaces. For the geometry, building on classical differential geometry, seminal work by David Huffman (1976) and subsequent developments (Duncan and Duncan, 1982; Fuchs and Tabachnikov, 1999; Kilian et al., 2008; Demaine et al., ) established a foundational theorem: the configuration of a curved-fold origami or, more generally, a developable surface is completely determined by the geometry of a single deformed curve on the surface—such as the crease in curved-fold origami. In other words, the distribution of curvature and torsion of the curve uniquely determines the shape of the entire surface. This theorem reveals an essential property of developable surfaces: despite being two-dimensional objects, their admissible deformations under isometric constraints are basically 1D, governed by the geometry of an underlying reference curve. For the statics, the differential geometry framework enables the modeling of panels as inextensible Kirchhoff plates with bending energy proportional to the mean curvature square. Following the geometry, the bending energy can be expressed as a functional of the curvature $\kappa$ and torsion $\tau$ of the reference curve, with the exact form given by the Wunderlich (1962) functional. Several approximated models have been developed for specific regimes, including the Sadowsky functional for narrow ribbons (Sadowsky, 1930) and Kirchhoff rod formulations for rod-like cross-sections. A comprehensive discussion of the applicability of these models—extensible versus inextensible, rod versus wide or narrow ribbons—is provided by Audoly and Neukirch (2021). In addition to the developable panels, the crease is typically modeled as a torsional spring, and the equilibrium configuration is obtained by minimizing the combined energy of the crease and panels, yielding shape equations governed by the curvature and torsion of the crease. This framework has been employed to analyze the deformation of folded circular strips (Dias et al., 2012), multi–curved-fold origami (Dias, 2012; Liu and James, 2024), and curved-fold origami with vertices (Wen et al., 2024).
@@
In contrast to the well-established geometric and static frameworks for developable surfaces and curved-fold origami, their dynamic behaviors and, in particular, the associated kinetic energy, remain poorly underdeveloped. Existing studies on dynamics are largely based on the dynamic Kirchhoff rod model for narrow and straight developable ribbons, according to previous works by Goriely et al. (2001), Dai et al. (2022) and Huang et al. (2024). Solving the Kirchhoff equations yields the motion of such structures. These models perform reasonably well for narrow ribbons (Yu et al., 2023; Shi et al., 2025). However, the Kirchhoff rod formulation does not enforce the isometric constraint of developable surfaces, leading to significant errors for highly twisted ribbons (Audoly and Neukirch, 2021), such as Möbius bands (Mahadevan and Keller, 1993; Starostin and van der Heijden, 2007). From the perspective of energy, this limitation arises because the line energy density of Kirchhoff rods scales as $\varepsilon_{crease}\in(A\kappa^{2}+B\tau^{2})$, which becomes inaccurate for configurations with large torsion (i.e. extremely curved case). In contrast, the Sadowsky functional exhibits the scaling $\varepsilon_{crease}\in(\kappa^{2}+\tau^{2})^{2}/\kappa^{2}$. To address this issue, Dias and Audoly (2014, 2015) developed a nonlinear rod model for narrow developable surfaces, employing constitutive laws derived from Sadowsky’s energy and enforcing the developability/isometric constraint. Despite these advances, current formulations remain limited to narrow ribbons and static settings. They do not provide a general dynamic theory applicable to developable surfaces of arbitrary width, which is required for large-scale applications such as space antennas.
@@
To this end, we begin with the kinematics to establish the correlations between the dynamic motions of the reference curve and the corresponding developable surface. Specifically, we analyze how a moving generatrix influences the surface velocity distribution and derive an explicit kinematic description of the velocity field regardless of the panel width and deformation. We thus derive the kinetic energy. Combined with the bending energy derived in statics, we then have the Lagrangian of the dynamic system. We find that the Lagrangian cannot be explicitly expressed as a functional of curvature $\kappa$ and torsion $\tau$, which prevents us from directly applying variational principles to derive the Euler-Lagrange equations of the system (as equations of $\kappa$ and $\tau$), as commonly done in static theories (Dias et al., 2012; Wen et al., 2024). Instead, leveraging the new Lagrangian, we introduce the panel-width effect, equivalent momentum and angular momentum into the nonlinear static rod model, thereby developing a nonlinear dynamic rod model for developable surfaces with arbitrary reference and deformed configurations. Solving the dynamical rod equations yields the motions of the reference curve and the developable surfaces under strict isometric constraint. We then extend our analysis to dynamic curve-fold origami. Composed of two developable surfaces joined by a curved fold, the kinematics of curve-fold origami is established similarly. By considering the dynamics of both panels and their geometric and mechanical correlations, we develop a nonlinear dynamical bi-rod model, which enables the study of the dynamics of a dynamic curve-fold origami. To validate the proposed dynamical theories, we conduct theoretical analyses on four representative cases: a rectangular ribbon, a curve-folded plate, a multi-curve-fold origami with vertices, and a folded ring origami. The theoretical results are validated by finite element analysis, demonstrating good agreement between the theory and simulation.
@@ 2.1 Preliminaries
to connect these two frames, and the $\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\}$ frame will not change its direction when moving along a generator. Thus the Darboux vectors only relate to the arc length $S$, following
@@ 2.2 Kinematics of developable surfaces
where $\mathbf{R}_{0}$ represents the position of the reference curve and $\mathbf{L}$ represents the generator in the reference configuration, which follows $\mathbf{L}=\eta\mathbf{E}_{1}+\mathbf{E}_{2}$ and $\mathbf{R}^{\prime}_{0}=\mathbf{E}_{1}$ as illustrated in Fig. 1(c). Equation (12) is thus transformed into
@@
$\dot{\mathbf{e}}_{i}=\bm{\Omega}\times\mathbf{e}_{i},$ (17)
Equation (16) can thus be written in a rather concise way
@@ 2.3 Lagrangian of developable surfaces: kinetic and elastic energy
where $D$ is the bending stiffness and $\varepsilon$ is the line density of elastic energy along the reference curve. Equation (20) indicates that $V$ is determined by the curvature of the reference curve and the generator distribution (or determined by the curvature and torsion of the reference curve).
@@
Equations (20) and (24) indicate that the Lagrangian, i.e., L=TVL=T-V, can be expressed as a one-dimensional integral along the reference curve. In statics analysis, we have L=VL=-V, which is a functional of $\kappa$ and $\tau$, and the Lagrangian can directly lead to the Euler-Lagrange equations in terms of $\kappa$ and $\tau$ (Dias et al., 2012; Wen et al., 2024). However, in dynamic analysis, the variables including $\mathbf{r}_{0}$ and $\bm{\Omega}$ cannot be explicitly expressed in terms of $\kappa$ and $\tau$, resulting in a complex Lagrangian that cannot directly lead to the governing equations. To address this issue, we draw inspiration from the equivalent rod model for static developable surfaces (Dias and Audoly, 2014, 2015) with narrow ribbons, and develop an equivalent model to handle the dynamical analysis for a general developable surface.
@@ 2.4 A 1D nonlinear dynamical model for general developable surfaces
As a conclusion of Section 2.3, the Lagrangian of a dynamic developable surface is a one-dimensional integral along the crease. Therefore, we can always establish a non-linear 1D dynamical model (similar to a dynamic rod model) with the same Lagrangian. The key point is that the two systems must have the same Lagrangian, which ensures that the motion of the reference curve on the developable surface matches that of the effective rod predicted by the 1D dynamic model. In this section, we derive explicit dynamical equations for the reference curve, similar to a nonlinear rod model. The new rod model will give the dynamics of the reference curve and thus the dynamics of the developable surface under isometric constraint. Since we use the Wunderlich functional (no narrow ribbon assumption), our model is an exact dimensional-reduced model that can predict the dynamics of a general developable surface accurately.
@@ 2.4.2 Dynamical equations
where $\gamma=\eta^{\prime}-(1+\eta^{2})\kappa_{g}$ and $m_{3}$ is a Lagrangian multiplier constrained by the geodesic curvature. The dynamic equation system (32) is one of the key contributions in this work. It employs the Wunderlinch functional for wide developable surfaces and extends the well-studied statics to the dynamic regime. The left-hand side of (32) considers the force/moment equilibrium along the effective rod, and the right-hand side contains the time evolution terms, resulting in the shape change with respect to time. The numerical scheme to solve the spatial-temporal evolution of (32) is given in B.
@@
Equations (32) and (33) characterize the full dynamical evolution of the reference curve, whose motion in turn determines the deformation of the entire developable surface.
@@ 2.5 Model validation: dynamics of a lifted rolled ribbon
In this section, we compare the results of our new model and the Kirchhoff rod model in the example shown in Fig. 3. As illustrated, a rectangular ribbon with aspect ratio $L/W=2$ is first rolled and lifted statically, and we study the dynamic process after the lifting load is released.
Figure 3: (a) A rectangular wide strip is rolled and lifted to the initial position. (b) The dynamic process after releasing the load. The color bar denotes the rotation angle around $z$ axis. (c) The quantitative results (curvature) predicted by our model (left) and the Kirchhoff rod model (right). Our model achieves better agreement with the FEM results.
@@
along with the boundary conditions $F=0$, $m=0$. Equation (36) is equivalent to the dynamical equations of the Kirchhoff rod model (Coleman et al., 1993; Goriely et al., 2001; Huang et al., 2024), which implies that the model established in this work can degenerate into existing static and dynamical models. In the Kirchhoff rod model (36), the curvature $\kappa=\kappa_{n}$ and torsion $\tau$ of the rod are decoupled, in contrast with the coupling of $\kappa$ and $\tau$ in our model. Figure 3 compares these two models to validate our new model’s ability to handle highly twisted panels with arbitrary width. The comparison between the theoretical results derived via different models along with the FEM results is shown in Fig. 3(c), illustrating that the Kirchhoff rod model has difficulties handling highly twisted wide panels, while the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces using our model.

### v3

@@ Abstract
Curve-fold origami, composed of developable panels joined along a curved crease, exhibits rich dynamic behaviors relevant to metamaterials and soft robotic systems. Despite multiple approximated models (shell model, Kirchhoff rod model, etc.), a comprehensive and exact dynamical theory for curve-fold origami remains absent, limiting the precise prediction of its dynamics, especially for those with wide panels. In this work, we develop an exact dimension-reduced theory that focuses on the dynamics of curve-fold origami, utilizing the intrinsic one-dimensional nature of developable surfaces. As is well known in differential geometry, a curve on a developable surface determines the local frame and the shape of the developable surface under isometric constraint, making the problem indeed one-dimensional. Starting from a single developable surface, we investigate the kinematics and kinetic energy of a moving developable surface. By overcoming the difficulty of describing the motion of local frames, we derive the exact velocity field of wide surfaces solely described by the motion of the reference curve, which leads to the kinetic energy of the entire surface. The elastic energy follows the Wunderlich functional applicable to wide panels. Owing to the one-dimensional feature, the Lagrangian of the system, composed of both kinetic and elastic energy, is a functional of the reference curve. Thus, we may variate the Lagrangian and derive a nonlinear dynamical theory for the reference curve, which comprises governing equations similar to the rod model but can precisely describe the motion of developable surfaces. The theory is validated consistently in both Lagrangian and Eulerian frameworks and is further extended to curved-fold origami modeled as a coupled bi-rod system. Utilizing our exact 1D model, we theoretically analyze the dynamical behaviors of various developable structures, revealing that the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces, which are validated by finite element analysis quantitatively. Our exact dimension-reduced dynamical theory captures the motions of developable surfaces and curve-fold origami (with wide panels) accurately at relatively low computational cost (since it is 1D), offering a potential contribution to the design of dynamical developable structures.
@@ 1 Introduction
Previous studies on the geometry and statics of curve-fold origami are under isometric assumptions, since the paper-like panels are unstretchable. Therefore, the panels will deform as developable surfaces. For the geometry, building on classical differential geometry, seminal work by David Huffman (1976) and subsequent developments (Duncan and Duncan, 1982; Fuchs and Tabachnikov, 1999; Kilian et al., 2008; Demaine et al., ) established a foundational theorem: the configuration of a curved-fold origami or, more generally, a developable surface is completely determined by the geometry of a single deformed curve on the surface—such as the crease in curved-fold origami. In other words, the distribution of curvature and torsion of the curve uniquely determines the shape of the entire surface. This theorem reveals an essential property of developable surfaces: despite being two-dimensional objects, their admissible deformations under isometric constraints are basically 1D, governed by the geometry of an underlying reference curve. For the statics, the differential geometry framework enables the modeling of panels as inextensible elastic plates with bending energy proportional to the mean curvature square. Following the geometry, the bending energy can be expressed as a functional of the curvature $\kappa$ and torsion $\tau$ of the reference curve, with the exact form given by the Wunderlich (1962) functional. Several approximated models have been developed for specific regimes, including the Sadowsky functional for narrow ribbons (Sadowsky, 1930) and Kirchhoff rod formulations for rod-like cross-sections. A comprehensive discussion of the applicability of these models—extensible versus inextensible, rod versus wide or narrow ribbons—is provided by Audoly and Neukirch (2021). In addition to the developable panels, the crease is typically modeled as a torsional spring, and the equilibrium configuration is obtained by minimizing the combined energy of the crease and panels, yielding shape equations governed by the curvature and torsion of the crease. This framework has been employed to analyze the deformation of folded circular strips (Dias et al., 2012), multi–curved-fold origami (Dias, 2012; Liu and James, 2024), and curved-fold origami with vertices (Wen et al., 2024).
@@
In contrast to the well-established geometric and static frameworks for developable surfaces and curved-fold origami, their dynamic behaviors and, in particular, the associated kinetic energy, remain poorly developed. Existing studies on dynamics are largely based on the dynamic Kirchhoff rod model for narrow and straight developable ribbons, according to previous works by Goriely et al. (2001), Dai et al. (2022) and Huang et al. (2024). Solving the Kirchhoff equations yields the motion of such structures. These models perform reasonably well for narrow ribbons (Yu et al., 2023; Shi et al., 2025). However, the Kirchhoff rod formulation does not enforce the isometric constraint of developable surfaces, leading to significant errors for highly twisted ribbons (Audoly and Neukirch, 2021), such as Möbius bands (Mahadevan and Keller, 1993; Starostin and van der Heijden, 2007). From the perspective of energy, this limitation arises because the line energy density of Kirchhoff rods scales as $\varepsilon_{crease}\in(A\kappa^{2}+B\tau^{2})$, which becomes inaccurate for configurations with large torsion (i.e., extremely curved case). In contrast, the Sadowsky functional exhibits the scaling $\varepsilon_{crease}\in(\kappa^{2}+\tau^{2})^{2}/\kappa^{2}$. To address this issue, Dias and Audoly (2014, 2015) developed a nonlinear rod model for narrow developable surfaces, employing constitutive laws derived from Sadowsky’s energy and enforcing the developability/isometric constraint. Despite these advances, current formulations remain limited to narrow ribbons and static settings. They do not provide a general dynamic theory applicable to developable surfaces of arbitrary width, which is required for large-scale applications such as space antennas.
@@
To this end, we begin with the kinematics to establish the correlations between the dynamic motions of the reference curve and the corresponding developable surface. Specifically, we analyze how a moving generatrix influences the surface velocity distribution and derive an explicit kinematic description of the velocity field regardless of the panel width and deformation. We thus derive the kinetic energy. Combined with the bending energy derived in statics, we then have the Lagrangian of the dynamic system. We find that the Lagrangian cannot be explicitly expressed as a functional of curvature $\kappa$ and torsion $\tau$, which prevents us from directly applying variational principles to derive the Euler-Lagrange equations of the system (as equations of $\kappa$ and $\tau$), as commonly done in static theories (Dias et al., 2012; Wen et al., 2024). Instead, leveraging the new Lagrangian, we introduce the panel-width effect, equivalent momentum and angular momentum into the nonlinear static rod model, thereby developing a nonlinear dynamic rod model for developable surfaces with arbitrary reference and deformed configurations. Solving the new dynamical rod equations yields the motions of the reference curve and the developable surfaces under strict isometric constraint. The comparison with classical Kirchhoff rod and plate models is presented to highlight the advantages of the proposed formulation. We then extend our analysis to dynamic curve-fold origami. Composed of two developable surfaces joined by a curved fold, the kinematics of curve-fold origami is established similarly. By considering the dynamics of both panels and their geometric and mechanical correlations, we develop a nonlinear dynamical bi-rod model, which enables the study of the dynamics of a dynamic curve-fold origami. To validate the proposed dynamical theories, we conduct theoretical analyses on four representative cases: a rectangular ribbon, a curve-folded plate, a multi-curve-fold origami with vertices, and a folded ring origami. The theoretical results are validated by finite element analysis, demonstrating good agreement between the theory and simulation.
@@ 2.1 Preliminaries
to connect these two frames, and the $\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\}$ frame will not change its direction when moving along a generator. Thus, the Darboux vectors only relate to the arc length $S$, following
@@ 2.2 Kinematics of developable surfaces
where $\mathbf{R}_{0}$ represents the position of the reference curve and $\mathbf{L}$ represents the generator in the reference configuration, which follows $\mathbf{L}=\eta\mathbf{E}_{1}+\mathbf{E}_{2}$ and $\mathbf{R}^{\prime}_{0}=\mathbf{E}_{1}$ as illustrated in Fig. 1(c). Eq. (12) is thus transformed into
@@
$\dot{\mathbf{e}}_{i}=\bm{\Omega}\times\mathbf{e}_{i}.$ (17)
Eq. (16) can thus be written in a rather concise way
@@ 2.3 Lagrangian of developable surfaces: kinetic and elastic energy
where $D$ is the bending stiffness and $\varepsilon$ is the line density of elastic energy along the reference curve. Eq. (20) indicates that $V$ is determined by the curvature of the reference curve and the generator distribution (or determined by the curvature and torsion of the reference curve).
@@
Eqs. (20) and (24) indicate that the Lagrangian, i.e., L=TVL=T-V, can be expressed as a one-dimensional integral along the reference curve. In statics analysis, we have L=VL=-V, which is a functional of $\kappa$ and $\tau$, and the Lagrangian can directly lead to the Euler-Lagrange equations in terms of $\kappa$ and $\tau$ (Dias et al., 2012; Wen et al., 2024). However, in dynamic analysis, the variables including $\mathbf{r}_{0}$ and $\bm{\Omega}$ cannot be explicitly expressed in terms of $\kappa$ and $\tau$, resulting in a complex Lagrangian that cannot directly lead to the governing equations. To address this issue, we draw inspiration from the equivalent rod model for static developable surfaces (Dias and Audoly, 2014, 2015) with narrow ribbons, and develop an equivalent model to handle the dynamical analysis for a general developable surface.
@@ 2.4 A 1D nonlinear dynamical model for general developable surfaces
As a conclusion of Section 2.3, the Lagrangian of a dynamic developable surface is a one-dimensional integral along the crease. Therefore, we can always establish a non-linear 1D dynamical model (similar to a dynamic rod model) with the same Lagrangian. The key point is that the two systems must have the same Lagrangian, which ensures that the motion of the reference curve on the developable surface matches that of the effective rod predicted by the 1D dynamic model. In this section, we derive explicit dynamical equations for the reference curve, similar to a nonlinear rod model. The new rod model will give the dynamics of the reference curve and thus the dynamics of the developable surface under isometric constraint. Since we use the Wunderlich functional (no narrow ribbon assumption), our model is an exact dimension-reduced model that can predict the dynamics of a general developable surface accurately.
@@ 2.4.2 Dynamical equations
where $\gamma=\eta^{\prime}-(1+\eta^{2})\kappa_{g}$ and $m_{3}$ is a Lagrangian multiplier constrained by the geodesic curvature. The dynamic equation system (32) is one of the key contributions in this work. It employs the Wunderlich functional for wide developable surfaces and extends the well-studied statics to the dynamic regime. The left-hand side of (32) considers the force/moment equilibrium along the effective rod, and the right-hand side contains the time evolution terms, resulting in the shape change with respect to time. The numerical scheme to solve the spatial-temporal evolution of (32) is given in B.
@@
Eqs. (32) and (33) characterize the full dynamical evolution of the reference curve, whose motion in turn determines the deformation of the entire developable surface.
@@ 2.5 Model validation: dynamics of a lifted rolled ribbon
In this section, we compare the results of our new model and the Kirchhoff rod model in the example shown in Fig. 3. As illustrated, a rectangular ribbon with aspect ratio $L/b=2$ is first rolled and split statically, and we study the dynamic process after the lifting load is released.
Figure 3: (a) A rectangular wide strip is rolled and lifted to the initial position. (b) The dynamic process after releasing the load. The color bar denotes the rotation angle around $z$ axis. (c) The quantitative results (nondimensionalized curvature $\kappa L$) predicted by our model (left) and the Kirchhoff rod model (right). Our model achieves better agreement with the FEM results. $T$ is the characteristic evolution time shown in (b).
@@
along with the boundary conditions $F=0$, $m=0$. Eq. (36) is equivalent to the dynamical equations of the Kirchhoff rod model (Coleman et al., 1993; Goriely et al., 2001; Huang et al., 2024), which implies that the model established in this work can degenerate into existing static and dynamical models. In the Kirchhoff rod model (36), the curvature $\kappa=\kappa_{n}$ and torsion $\tau$ of the rod are decoupled, in contrast with the coupling of $\kappa$ and $\tau$ in our model. Fig. 3 compares these two models to validate our new model’s ability to handle highly twisted panels with arbitrary width. The comparison between the theoretical results derived via different models along with the FEM results is shown in Fig. 3(c), illustrating that the Kirchhoff rod model has difficulties handling highly twisted wide panels, while the coupling of curvature and torsion along with the motion of local frames in our theory leads to the accurate modeling of arbitrarily deformed developable surfaces using our model.
2.6 Model validation: transient response of an elastic plate
In the field of thin sheets and plates, the Kirchhoff–Love plate model is also a classic and commonly used model. In this section, we compare our new model and the Kirchhoff-Love plate model in the example shown in Fig. 4. A plate (length: $L$, width: $b$, thickness: $h$) with simply supported sliding ends is subjected to an instantaneous velocity field following
$\dot{w}(x,y)=0.3\frac{\pi^{2}}{L^{2}}\sqrt{\frac{D}{\rho h}}\sin\frac{\pi x}{L}$ (37)
where $w$ is the vertical displacement. The plate consequently begins to vibrate, as illustrated in Fig. 4(a). Following the $y$ invariance, the dynamic equations Eq. (32) are thus simplified into
$F_{w}\cos\varphi-F_{x}\sin\varphi+D\varphi^{\prime\prime}=0,$ (38)
$F_{x}=\rho h\ddot{x},\quad F_{w}=\rho h\ddot{w},$
$x^{\prime}=\cos\varphi,\quad w^{\prime}=\sin\varphi,$
along with the boundary conditions
$F_{x}=0,\quad\varphi^{\prime\prime}=0,\quad w=0,$ (39)
at both ends. Solving Eqs. (38) and (39) yields the motion of the plate.
Figure 4: (a) The vibration of an elastic plate with an initial velocity field. (b) The configuration (in the xwxw plane) predicted by our model and the Kirchhoff-Love plate model. Our model achieves better agreement with the FEM results. The length $L$ is used to nondimensionalize the geometry.
In Kirchhoff-Love plate theory, the dynamic equation follows (Kirchhoff, 1850)
$Dw^{\prime\prime\prime\prime}+\rho h\ddot{w}=0.$ (40)
With the boundary conditions Eq. (39), the explicit solution of Eq. (40) is
$w=A\sin\frac{\pi S}{L}\sin(\frac{\pi^{2}}{L^{2}}\sqrt{\frac{D}{\rho h}}t),$ (41)
where A is the amplitude determined by the initial velocity field.
To validate the two theoretical models, we derive the configuration at time $T$ (defined as the time when the plate first reaches zero velocity) and $0.5T$ from both Kirchhoff-Love plate theory and the new theory. As illustrated in Fig. 4(b), the solution of the new theory aligns better with finite element results, indicating that the new theory can handle large curvature problems more accurately.
As a conclusion of Sections 2.5 and 2.6, we summarize the assumptions, advantages and limitations of the Kichhoff-rod model, Kirchhoff-Love plate model and our new model in the following chart.